ADNAN MUHAMMED

Grade 12,

At what temperature the most probable velocity of oxygen gas is equal to the r.m.s velocity of ozone at t degree celcius

At what temperature the most probable velocity of oxygen gas is equal to the r.m.s velocity of ozone at t degree celcius

Grade:11

1 Answers

Onkar Joshi
6 years ago
Temperature of ozone=t°C=( 273+t)Krms velocity of O3 is given by=√3RT/M=√3R(273+t)/0.048Most probable speed of O2=√2RT/M=√2RT/0.032√3R(273+t)/0.048=√2RT/0.032T=273+t K= t°CAt t°C the MPs of O2= rms of O3